What is the probability that a 5 card poker hand has 2 queens, given that all 5 cards are red cards?
My impression is this is an intersection problem { P(X and Y) = P(X) + P(Y) - P(X)*P(Y) } ?

Answer by stanbon(75887)

What is the probability that a 5 card poker hand has 2 queens, given that all 5 cards are red cards?


# of ways to succeed: 2C2*24C3 = 2024
-----------
# of possible 5 red-card hands: 26C5 = 65780
-------------
P(condition met) = 2024/65780 = 0.0308
-----
Cheers,
Stan H.